### v1

# Large Language Models as Urban Residents: An LLM Agent Framework for Personal Mobility Generation

Abstract
This paper introduces a novel approach using Large Language Models (LLMs) integrated into an agent framework for flexible and efficient personal mobility generation. LLMs overcome the limitations of previous models by efficiently processing semantic data and offering versatility in modeling various tasks. Our approach addresses the critical need to align LLMs with real-world urban mobility data, focusing on three research questions: aligning LLMs with rich activity data, developing reliable activity generation strategies, and exploring LLM applications in urban mobility. The key technical contribution is a novel LLM agent framework that accounts for individual activity patterns and motivations, including a self-consistency approach to align LLMs with real-world activity data and a retrieval-augmented strategy for interpretable activity generation. In experimental studies, comprehensive validation is performed using real-world data. This research marks the pioneering work of designing an LLM agent framework for activity generation based on real-world human activity data, offering a promising tool for urban mobility analysis.
1 Introduction
Under the 2030 Sustainable Development Agenda, the United Nations emphasizes the crucial role of sustainable cities and human settlements. The study of personal mobility like individual activity trajectory can provide valuable semantic insights into urban mobility and support this objective. Specifically, modeling personal mobility opens up numerous applications for building a sustainable community, including proactive traffic management and the design of comprehensive urban development strategies Batty et al. (2012); Batty (2013); Zheng (2015). In particular, generating reliable activity trajectories can be a promising and efficient way to exploit individual activity data. On one hand, learning to generate activity trajectory leads to a thorough understanding of activity patterns, enabling flexible simulation of urban mobility. On the other hand, while individual activity trajectory data is abundant due to advances in telecommunications, its practical use is often limited by privacy concerns. In this sense, generated data can provide a viable alternative that offers a balance between utility and privacy.
Figure 1: Personal mobility generation with an LLM agent.
While advanced data-driven learning-based method offer various solutions to generate synthetic individual trajectoriesHuang et al. (2019); Zhang et al. (2020); Feng et al. (2020); Choi et al. (2021), the generated data only imitates real-world data from the data distribution perspective rather than semantics, rendering them less effective in simulating or interpretation activities in novel or unforeseen scenarios with a significantly different distribution (e.g., a pandemic). Thus, in this study, to explore a more intelligent and efficient activity generation, we propose to establish an agent framework by exploiting the emerging intelligence of large language models (LLMs), as illustrated by Fig. 1. LLMs have gained widespread usage in furthering our understanding of humans and society in a multitude of disciplines, such as political science Argyle et al. (2023) and economy Aher et al. (2022), and thus have been employed as agents in a variety of social simulations Xi et al. (2023); Wu et al. (2023); Gao et al. (2023). More specifically, LLMs present two significant advantages over previous models when applied to activity trajectory generation:
  • Semantic Interpretability. Unlike previous models, which have predominantly depended on structured data (e.g., GPS coordinates-based trajectory data) for both calibration and simulation Jiang et al. (2016); Pappalardo and Simini (2018), LLMs exhibit proficiency in interpreting semantic data (e.g., activity trajectory data). This advantage significantly broadens the scope for incorporating a diverse array of data sources into generation processes, thereby enhancing the models’ ability to understand and interact with complex, real-world scenarios in a more nuanced and effective manner.
  • Model Versatility. Although other data-driven methods manage to learn such dynamic activity patterns for generation, their capacity is limited for generation under unseen scenarios. On the contrary, LLMs have shown remarkable versatility in dealing with unseen tasks, especially the ability to reason and decide based on available information OpenAI (2022). This competence enables LLMs to offer a diverse and rational array of choices, making it a promising and flexible approach for modeling personal mobility patterns.
Despite these benefits, ensuring that LLM aligns effectively with real-world situations continues to be a significant challenge Xi et al. (2023). This alignment is particularly crucial in the context of urban mobility, where the precision and dependability of LLM outputs are essential for the efficacy of any urban management derived from them. In this study, our aim is to address this challenge by investigating the following research questions: RQ 1: How can LLM be effectively aligned with semantically rich data about daily individual activities? RQ 2: What are the effective strategies for achieving reliable and meaningful activity generation using LLM? RQ 3: What are the potential applications of LLM in enhancing urban mobility?
Our methodology revolves around two phases: (1) self-consistent activity pattern identification and (2) motivation-driven activity generation. In Phase 1, we extract and evaluate activity patterns from historical data and semantic analysis, using LLM agents to generate diverse and personalized activity scenarios. This process involves identifying habitual activity patterns, which are key factors influencing individual mobility. In Phase 2, we develop two interpretable retrieval-augmented strategies, along with the patterns identified in Phase 1, to instruct LLM agents to work out summarized motivations, such as evolving interests or situational needs. The patterns and motivations faciliate the agents to generate daily activity trajectories. The generation benefits from the LLM’s capability to process extensive datasets and semantic richness, allowing for a nuanced simulation of personal mobility with considerations for habitual behaviors, situational motivations, and temporal dynamics.
We evaluate the proposed framework using GPT-3.5 APIs over a personal activity trajectory dataset of Tokyo. The results demonstrate the capability of our framework to align LLM agents with semantically rich data for generating individual daily activities, highlighting that our framework excels in temporal aspects of personal mobility generation, with spatial-temporal metrics showing its strength in capturing the essence of daily routines over mere location accuracy. The comparison with baselines like Markov models, LSTM, and DeepMove Feng et al. (2018), underscores the advanced generative performance of our framework. Moreover, the application of the framework in simulating urban mobility under specific contexts, such as the pandemic scenario, reveals its potential to adapt to external factors and generate realistic activity patterns.
To the best of our knowledge, this study is one of the pioneering works in developing an LLM agent framework for generating activity trajectory based on real-world data. We summarize our contributions as follows: (1) We introduce a novel LLM agent framework for urban mobility generation featuring semantic richness. (2) Our framework introduces a self-consistency evaluation to ensure that the output of the LLM aligns closely with real-world data on daily activities. (3) To generate daily activity trajectories, our framework integrates activity patterns with summarized motivations, with two interpretable retrieval-augmented strategies aimed at producing reliable activity trajectories. (4) By using real-world personal activity data, we validate the effectiveness of our framework and explore its utility in urban mobility analysis.
2 Related Work
2.1 Personal Mobility Generation
Activity trajectory generation offers a valuable perspective for understanding urban mobility. Benefiting from the advanced communication technologies, recent research can explore urban mobility from massive geotagged data. Based on vast call detailed records, Jiang et al.Jiang et al. (2016) built a mechanistic modeling framework to generate individual activities in high spatial-temporal resolutions, but it is limited to three activity types: work, home, and others. Pappalardo and SiminiPappalardo and Simini (2018) enhanced the Exploration and Preferential Return (EPR) model for activity generation, which employs Markov modeling to estimate the probability of individuals visiting specific locations. Besides, deep learning has become a robust tool for modeling the complex dynamics of traffic  Liu et al. (2018); Huang et al. (2019); Zhang et al. (2020); Feng et al. (2020); Choi et al. (2021); Luca et al. (2021). The primary challenge involves overcoming data-related obstacles such as randomness, sparsity, and irregular patterns Feng et al. (2018); Yuan et al. (2023, 2022); Long et al. (2023). For example, Feng et al.Feng et al. (2018) proposed attentional recurrent networks to handle personal preference and transition regularities. Yuan et al.Yuan et al. (2022) leveraged deep learning combined with neural differential equations to address the challenges of randomness and sparsity inherent in irregularly sampled activities for activity trajectory generation.
Figure 2: Illustration of LLMob, the proposed LLM agent framework for personal Mobility generation.
2.2 LLM Agents
How to take advantage of LLMs for real-world tasks has become an attractive research topic since the LLMs present human-like intelligence. While some studies Wei et al. (2022); Wang et al. (2022) focus on exploiting their linguistic capabilities, exploring how to treat LLM as autonomous agents in specific scenarios can lead to more diverse and promising applications Wang et al. (2023); Xi et al. (2023); Wu et al. (2023); Gao et al. (2023), hence to address the issues of reliance on parameter settings and lack of adaptability in traditional agent-based simulations using rules or reinforcement learning Grimm et al. (2005); Henrich et al. (2010); Ale Ebrahim Dehkordi et al. (2023). For instance, Park et al.Park et al. (2023) first established an LLM-powered agent framework to simulate human behavior in an interactive scenario. This innovative approach demonstrates the potential of LLM to model complex social interactions and decision-making processes. Extending beyond the realm of social sciences, Mao et al.Mao et al. (2023) have adeptly utilized LLM to generate driving trajectories in motion planning tasks. In the field of natural sciences, Williams et al.Williams et al. (2023) have innovatively integrated LLM with epidemic models to simulate the spread of diseases. Furthermore, in financial research, the application of LLM in macroeconomic modeling has been explored, providing new insights into financial markets and economies Li et al. (2023). These varied applications highlight the versatility and potential of LLM to understand and model various real-world dynamics.
3 Problem Description
We study the generation of individual daily activity trajectories, each representing an individual’s activities for the whole day. We focus on the urban context, where the activities of each individual are represented as a time-ordered sequence of location choices (e.g., location-based activity) Luca et al. (2021). This sequence, $\{(l_{0},t_{0}),(l_{1},t_{1}),\ldots,(l_{n},t_{n})\}$, represents the activity trajectory of an individual (e.g., individual POI trajectories), where $(l,t)$ denotes the individual’s location $l$ at time $t$.
4 Methodology
By modeling individuals within an urban environment as LLM agents, we present LLMob, an LLM Agent Framework for Personal Mobility Generation, as illustrated in Fig. 2. LLMob is based on the assumption that an individual’s activities are primarily influenced by two principal factors: habitual activity patterns and current motivations. Habitual activity patterns, which reflect typical movement behaviors and preferences, signifying regular travel and location choices Song et al. (2010). Motivations pertain to dynamic and situational factors that influence an individual’s decisions at any given moment, such as specific needs or external influences during a particular time frame. This consideration is vital for capturing and forecasting short-term shifts in mobility patterns Aher et al. (2022). Moreover, by formulating prompts that assume specific events of concern, this framework allows us to observe the agents’ activities and responses in a variety of situations.
To construct a pipeline for activity trajectory generation, we design LLM agents with action, memory, and planning, in line with Weng (2023); Wang et al. (2023) except that external tools are not involved. Action specifies how an agent interacts with the environment and makes decisions. In LLMob, the environment contains the information collected from real-world data, and the agents act by identifying habitual activity patterns and generating trajectories. Memory includes past actions that need to be prompted to the LLM to invoke the next action. In LLMob, memory refers to the active patterns output by the agents. Planning formulates a plan or reflects over past actions to handle complex tasks. In LLMob, planning refers to the generation of summarized motivations, which are used to facilitate trajectory generation. Furthermore, we specify personas to the agents to simulate the diversity of real-world individuals Salewski et al. (2023).
4.1 Self-Consistent Activity Pattern Identification
Phase 1 of LLMob focuses on identifying activity patterns from historical data. We leverage the extracted activity patterns as essential prior knowledge for the generation of daily activities. To effectively identify activity patterns from historical data, we introduce the following two steps.
4.1.1 Pattern Extraction from Semantics and Historical Data
The first step will use LLM to derive activity patterns based on activity trajectory data (e.g. individual check-in data). As illustrated in the left panel of Fig. 2, this scheme consists of three aspects: For each person, we start by utilizing the LLM to create a variety of personas, providing the inspiring foundation for subsequent activity pattern generation. This approach also encourages the diversity of the generated activity patterns, as each persona acts as a unique prior for the generation process (e.g., Noulas et al.Noulas et al. (2011) demonstrated the significance of user clustering from activity trajectory data in producing meaningful distinctions). Meanwhile, we perform data preprocessing to extract key information from the extensive historical data. This involves identifying usual commuting distances, pinpointing typical start and end times and locations of daily trips, and concluding the most frequently visited locations of the person. It is important to note that these pieces of information are widely recognized as critical features in mobility analysis Jiang et al. (2016). Following this, both semantic elements with historical data are combined in the prompts, requiring the LLM to summarize the activity patterns for this person. By doing this, we set up a streamline to effectively bridge the gap between semantic persona characteristics and concrete historical activity trajectory data, which allows for a more personalized and interpretable representation of individual activities at one day.
4.1.2 Pattern Evaluation with Self-Consistency
As introduced before, for each person, multiple candidate patterns are generated according to the persona and historical data. Following this, the subsequent step involves assessing the consistency of these generated activity patterns to identify the most plausible one. We implement a scoring mechanism to gauge the alignment of these patterns with historical data. To achieve this objective, we define a scoring function for each candidate pattern $cp$ within the set $\mathcal{CP}$. This function evaluates $cp$ against two distinct sets of activity trajectories: the specific activity trajectories $\mathcal{T}_{i}$ of a targeted user $i$ and the sampled activity trajectories from other users $\mathcal{T}_{\sim i}$:
$score_{cp}=\sum_{t\in\mathcal{T}_{i}}r_{t}-\sum_{t^{\prime}\in\mathcal{T}_{\sim i}}r_{t^{\prime}},$ (1)
where we design an evaluation prompt to ask the LLM to generate rating scores $r_{t}$ and $r_{t^{\prime}}$. Specifically, the LLM is prompted to assess the degree of preference for a given trajectory based on the candidate pattern. Ideally, the LLM should assign a higher $r_{t}$ for data from the targeted user and a lower $r_{t^{\prime}}$ for data from other users. This scheme requires self-consistency in the pattern: the activity pattern derived from the activity trajectory data should be consistent with the data during the evaluation. The algorithm for this step is summarized as shown in Alg. 1.
Input : Number of personas $C$, activity trajectories $\mathcal{T}$
Output : Best activity pattern $best\_pattern$
1 Initialize: Empty candidate pattern set $\mathcal{CP}$;
2 Formulate initial pattern summarization prompt $p_{1}$ using prior information extracted from $\mathcal{T}$;
3 for $i=1$ to $C$ do
4       Formulate $p_{1}$ with prior information and persona $i$;
5       $cp\leftarrow\text{LLM}(p_{1})$;
6       $\mathcal{CP}\leftarrow\mathcal{CP}\cup\{\,cp\,\}$;
7
8Initialize a score dictionary $\mathcal{S}$ to store pattern scores;
9 foreach candidate pattern $cp$ in $\mathcal{CP}$ do
10       Initialize $score_{cp}=0$;
11       foreach activity trajectory $t$ in $\mathcal{T}$ do
12             Formulate evaluation prompt $p_{2}$ for $cp$ and $t$;
13             $score_{cp}\leftarrow score_{cp}+\text{LLM}(p_{2})$;
14      $\mathcal{S}[cp]\leftarrow score_{cp}$;
15
[…]
return $best\_pattern$
Algorithm 1 Self-Consistent Activity Pattern Identification
4.2 Motivation-Driven Activity Generation
In Phase 2 of LLMob, we focus on the retrieval of motivation and the integration of motivation and activity patterns for individual activity trajectory generation. Retrieval augmentation has been identified as a crucial factor in boosting the performance of LLM Xu et al. (2023). This enhancement provides additional information that aids LLM in more effectively responding to queries. While previous studies on activity generation mainly overlook the critical factors of macro temporal information (i.e., date) or specific scenarios (i.e., hash weather) Yuan et al. (2022), we propose a more sophisticated activity generation which accounts for various conditions by taking advantage of the human-like intelligence of LLM. For instance, the activity trajectory at date $d$ can be inferred given the motivation of this date and the normal activity pattern as:
$\mathcal{T}_{d}=LLM(\mathcal{M}otivation,\mathcal{P}attern).$ (2)
This activity generation scheme instructs the LLM to simulate a designated individual according to a given activity pattern, and then meticulously generate an activity trajectory in accordance with the daily motivation. To obtain insightful and reliable motivations for the generation, two retrieval schemes are proposed.
4.2.1 Evolving-based Motivation Retrieval
This scheme is related to the intuitive principle that an individual’s motivation on any given day is influenced by her interests and priorities in preceding days Park et al. (2023). Guided by this understanding that new motivations often evolve from existing ones, our approach harnesses the intelligence of LLM to understand the behavior of daily activities and the underlying motivations. As illustrated in Fig. 3, for a specific day $d$ for which we aim to generate the activity trajectory, as well as considering the activities of the past few days, we prompt the LLM to act as an individual based on the pattern identified in Section 4.1 and summarize the motivation behind these activities. Leveraging these identified past motivations, the LLM is further prompted to infer potential motivations for the subsequent day.
Figure 3: Evolving-based motivation retrieval.
4.2.2 Learning-based Motivation Retrieval
In this scheme, we hypothesize that individuals tend to establish routines in their daily activities, guided by consistent motivations even if the specific locations may vary. For example, if someone frequently visits a burger shop on weekday mornings, this behavior might suggest a motivation for a quick breakfast. Based on this, it is plausible to predict that the same individual might choose a different fast food restaurant in the future, motivated by a similar desire for convenience and speed during their morning meal. We introduce a learning-based scheme to retrieve motivation from historical data. For each new date on which to plan activities, the only information available is the date itself. To use this clue for planning, we first formulate a relative temporal feature $\boldsymbol{z}_{{d_{c},d_{p}}}$ between a past date $d_{p}$ and the current date $d_{c}$. This feature captures various aspects, such as the gap between these two dates and whether they belong to the same month. Utilizing this setting, we train a score approximator $f_{\theta}(\boldsymbol{z}_{{d_{c},d_{p}}})$ to evaluate the similarity between any two dates. Notably, due to the lack of supervised signals, we employ unsupervised learning to train $f_{\theta}(\cdot)$. Particularly, a learning scheme based on contrastive learning Chen et al. (2020) is established. For each trajectory of a user, we can scan her other trajectories and identify similar (positive) and dissimilar (negative) dates according to a predefined similarity score. This similarity score is calculated between two activity trajectories $\mathcal{T}_{d_{a}}$ and $\mathcal{T}_{d_{b}}$ as:
$sim_{d_{a},d_{b}}=\sum_{t=1}^{N_{d}}\mathbf{1}_{(\mathcal{T}_{d_{a}}(t)=\mathcal{T}_{d_{b}}(t))}\text{ if }|\mathcal{T}_{d_{a}}|>t\text{ and }|\mathcal{T}_{d_{b}}|>t,$ (3)
where $N_{d}$ is the total number of time intervals (e.g., 10 min) in one day. $\mathcal{T}_{d_{a}}(t)$ indicates the $t$th visiting location recorded in trajectory $\mathcal{T}_{d_{a}}$. Intuitively, there should be more shared locations in the similar trajectory pair. Thereafter, positive pair is characterized by the highest similarity score, indicative of a greater degree of resemblance between the trajectories. Conversely, negative pairs are marked by low similarity scores, reflecting a lesser degree of commonality. After obtaining the training dataset from these positive and negative pairs, we can train a model to approximate the similarity score between any two dates through contrastive learning. Specifically, the procedure involves the following steps:
  1. 1.
    For each date $d$, generate one positive pair $(d,d^{+})$ and $k$ negative pairs ($d,d^{-}_{1}$), …, ($d,d^{-}_{k}$) based on the similarity score and compute $\boldsymbol{z}_{d,d^{+}}$, $\boldsymbol{z}_{d,d^{-}_{1}}$, …, $\boldsymbol{z}_{d,d^{-}_{k}}$.
  2. 2.
    Forward the positive and negative pairs to $f_{\theta}(\cdot)$ to form:
    $\text{logits}=\left[f_{\theta}(\boldsymbol{z}_{d,d^{+}}),f_{\theta}(\boldsymbol{z}_{d,d^{-}_{1}}),...,f_{\theta}(\boldsymbol{z}_{d,d^{-}_{k}})\right].$ (4)
  3. 3.
    Adopt InfoNCE Oord et al. (2018) as the contrastive loss function:
    $\mathcal{L}(\theta)=\sum_{n=1}^{N}-\log\left(\frac{e^{\text{logits}_{i}}}{\sum^{k+1}_{j=1}e^{\text{logits}_{j}}}\right)_{n},$ (5)
    where $N$ is the batch size of the samples, $i$ indicates the index of the positive pair.
Upon training a similarity score approximation model, it can be applied to access the similarity between any given query date and historical dates. This enables us to retrieve the most similar historical data, which allows the LLM to generate a summary of the motivations prevalent at that time. By doing so, we can extrapolate a motivation relevant to the query date, providing a basis for the LLM to generate a new activity trajectory.
5 Experiment
5.1 Experimental Setup
Dataset. We investigate and validate LLMob over a personal activity trajectory dataset from Tokyo. This data set is obtained through the Twitter API and the Foursquare API, which covers data from January 2019 to December 2022.
Table 1: Sample of the personal activity data.
UserID Latitude Longitude Location Name Category Time 12199425 35.508 139.615 Convenience Store Shop & Service 8:00 12199425 35.509 139.618 Seafood Restaurant Food & Service 8:30
The time frame of this data set is insightful as it captures typical daily life prior to the COVID-19 pandemic (i.e., normal period) and subsequent alterations during the pandemic (i.e., abnormal period). To facilitate an efficient and detailed analysis for different periods, we randomly choose 100 users to model their individual activity trajectory at a 10-minute interval according to the number of available trajectories (sample given in Table 1). We utilize the category classification in Foursquare to determine the activity category for each location. We use 10 candidate personas as a prior for subsequent pattern generation, as shown in Table 2, which adequately capture a diverse range of activity patterns within the data of this study. For the application to other datasets, this style of candidate patterns can be easily initialized using an LLM.
Table 2: Suggested personas and corresponding detailed description.
Student: typically travel to and from educational institutions at similar times.
Teacher: typically travel to and from educational institutions at similar times.
Office worker: have a fixed morning and evening commute,
often heading to office districts or commercial centers.
Visitor: tend to travel throughout the day,
often visit attractions, dining areas, and shopping districts.
Night shift worker: might travel outside of standard business hours,
often later in the evening or at night.
Remote worker: may have non-standard travel patterns,
often visit coworking spaces or cafes at various times.
Service industry worker: tend to travel throughout the day, often visit attractions,
dining areas, and shopping districts.
Public service official: often work in shifts,
leading to varied travel times throughout the day and night.
Fitness enthusiast: often travel early in the morning, in the evening,
or on weekends to fitness centers or parks.
Retail employee: typically travel to and from educational institutions at similar times.
Metrics. To evaluate the generation performance, various features related to individual activity are used as metrics: (1) Step Distance (SD)Yuan et al. (2022): This metric evaluates the spatial pattern of an individual’s activities by measuring the distance between two consecutive locations in a trajectory. (2) Step Interval (SI)Yuan et al. (2022): This metric evaluates the temporal pattern of an individual’s activities by measuring the time interval between two successive locations on an individual’s trajectory. (3) ST-ACT: This metric captures and examines the spatial-temporal distribution of activities for each trajectory, focusing on the patterns of individual activities characterized by activity type and timing (e.g., activity type, time). It provides insight into how activities are distributed over space and time. (4) ST-LOC: This metric provides a granular perspective on the generated activities by assessing the spatial-temporal distribution of visited locations within each trajectory, including geographical coordinates and timestamps. It enables a detailed analysis of where and when activities occur.
After extracting the above features from both the generated and real activity trajectory data, Jensen-Shannon divergence (JSD) is employed to quantify the discrepancy between the generated and the actual ones. Lower JSD is preferred.
Baselines. The following models are implemented to serve as baselines: (1) Markov-based mechanic model (MM) Pappalardo and Simini (2018), (2) LSTM-based prediction model (LSTM), (3) Attentional Recurrent Network-based prediction model (DeepMove) Feng et al. (2018).
[…]
5.2 Experimental Results and Analysis
Generative Performance Validation (RQ1, RQ2). The performance evaluation involves analyzing generation results in three distinct settings: (1) Generating normal trajectories based on normal historical trajectories, specifically using data from 2019, a period unaffected by the pandemic. (2) Generating abnormal trajectories based on abnormal historical trajectories, utilizing data from 2020, a year marked by the pandemic. (3) Generating abnormal trajectories based on normal historical trajectories, using data of 2019 and 2021. Through these settings, the efficiency of aligning real-world data, along with the effectiveness of identifying patterns and motivations, can be validated. Furthermore, we expect to examine the models’ generalization capabilities across varying data situations. The results of these evaluations are detailed in the metrics reported in Tables 3 – 5. Through the comparison, it can be observed that although LLMob may not excel in replicating spatial features precisely (such as SD), it demonstrates superior performance in handling temporal aspects. Notably, when considering spatial-temporal features (like ST-ACT and ST-LOC), LLMob’s performance is markedly better. Specifically, we suggest that the pronounced advantage in terms of ST-ACT can be attributed to the LLM agents’ tendency to accurately replicate the motivation behind individual activity behaviors. For instance, they may recognize patterns like a person’s habits to have breakfast in the morning, without being restricted to a specific restaurant. This phenomenon highlights the enhanced semantic understanding capabilities of the LLM agents.
Table 3: Performance (JSD) of normal trajectory generation based on normal historical data.
Models SD SI ST-ACT ST-LOC
MM 0.018 0.276 0.644 0.681
LSTM 0.017 0.271 0.585 0.652
DeepMove 0.008 0.153 0.534 0.623
LLMob-E 0.053 0.046 0.125 0.559
LLMob-L 0.049 0.054 0.136 0.570
LLMob-L w/o $\mathcal{P}$ 0.061 0.080 0.270 0.600
LLMob-E w/o $\mathcal{P}$ 0.055 0.069 0.223 0.530
[…]
Table 4: Performance (JSD) of abnormal trajectory generation based on abnormal historical data.
Models SD SI ST-ACT ST-LOC
MM 0.041 0.300 0.629 0.682
LSTM 0.016 0.286 0.563 0.655
DeepMove 0.011 0.173 0.548 0.668
LLMob-E 0.056 0.043 0.127 0.615
LLMob-L 0.057 0.051 0.124 0.609
LLMob-L w/o $\mathcal{P}$ 0.072 0.081 0.286 0.641
LLMob-E w/o $\mathcal{P}$ 0.059 0.081 0.252 0.573
[…]
Table 5: Performance (JSD) of abnormal trajectory generation based on normal historical data.
Models SD SI ST-ACT ST-LOC
MM 0.039 0.307 0.644 0.681
LSTM 0.035 0.282 0.585 0.653
DeepMove 0.013 0.173 0.534 0.623
LLMob-E 0.072 0.056 0.117 0.536
LLMob-L 0.064 0.051 0.124 0.531
LLMob-L w/o $\mathcal{P}$ 0.073 0.091 0.248 0.580
LLMob-E w/o $\mathcal{P}$ 0.065 0.079 0.209 0.561
[…]
Figure 4: Daily activity frequencies comparison.
Figure 5: Evolving of accumulative counts, Arts & Entertainment.
Figure 6: Evolving of accumulative counts, Professional & Other Places.
Explore Real-World Application (RQ3). We are interested in how LLMob can elevate the social benefits, particularly in the context of urban mobility. To this end, we propose an example of leveraging the flexibility and intelligence of the LLM in understanding semantic information. In particular, we enhance the original setup by incorporating an additional prompt to provide a context for the LLM agent, enabling it to plan activities during specific circumstances. For example, a “pandemic” prompt is as follows: Now it is the pandemic period. The government has asked residents to postpone travel and events and to telecommute as much as possible.
By integrating this prompt, we introduce an external factor – the pandemic situation and the government’s response – into LLMob. In this way, we can observe and analyze the impact of external elements, such as the ongoing pandemic and the government’s measures, on urban mobility and related social dynamics. We use the activity trajectory data during the pandemic (e.g., 2021) as ground truth. As shown in Fig. 4, “LLMob-L” serves as the activity trajectory generator, and the one augmented with the pandemic prompt is denoted as LLMob-L with “Pandemic” Prompt. The influence of the external factor on the simulation is clear: there is a significant decrease in activity frequency when the prompt, which semantically discourages activities prone to spreading the disease, is applied. Additionally, from a spatial-temporal perspective, from spatial-temporal perspective, two major activities (e.g., Arts & Entertainment and Professional & Other Places) are selected to observe the behavior, as shown in Figs. 5 and  6. These activities are particularly insightful as they encapsulate the impact of the pandemic on the work-life balance and daily routines of residents.
Specifically, “LLMob-L with Pandemic Prompt” reproduces a more realistic spatial-temporal activity pattern. This enhanced realism in the generation is attributed to the integration of prior knowledge about the pandemic’s effects and governmental responses, allowing the LLM agent to behave in a manner that aligns with actual behavioral adaptations. For instance, the reduction in Arts & Entertainment activities reflects the closure of venues and social distancing guidelines, while changes in Professional & Other Places activities indicate shifts toward remote work and the transformation of professional environments. Intuitively, prompting LLM agents to generate activities based on various priors shows great potential in real-world applications. The utility of such a conditioned generative approach, coupled with the reliable generated results, can significantly alleviate the workload of urban managers. We suggest that this kind of workflow can simplify the analysis of urban dynamics and aid in assessing the potential impact of urban policies.
6 Conclusion
Contribution. This study is believed to be the first urban mobility simulation empowered by LLM on real-world data. Our innovative framework leverages activity patterns and motivations to direct LLM agents in emulating urban residents, facilitating the generation of interpretable and efficient individual activity trajectory. Extensive experimental studies based on real-world data are conducted to validate the proposed framework. Further exploration demonstrated the promising capabilities of LLM to improve urban mobility management.
Social Impact. Leveraging artificial intelligence to enhance societal benefits is increasingly promising, especially with the advent of high-capacity models such as LLMs. This study explores one of the potential avenues for applications using LLMs as a reliable agent to simulate specific scenarios to assess the effects of external factors, such as government policies. The introduced framework offers a flexible approach to enhance the reliability and efficiency of LLMs in simulating urban mobility.
Acknowledgements
This work is supported by JSPS Kakenhi 22H03903, 23H03406, 23K17456, and JST CREST JPMJCR22M2 and SPRING JPMJSP2108.
Ethical Statement
Despite the use of GPT for polishing the paper, we are responsible for all content. The data is anoynimized. To the best of our knowledge, we find no ethical issues.
Appendix
Appendix A Prompts
Pattern generation Context: Act as a $<$INPUT 0$>$ in an urban neighborhood, $<$INPUT 1$>$ $<$INPUT 2$>$. There are also locations you sometimes visit: $<$INPUT 3$>$. Instructions: Reflecting on the context given, I’d like you to summarize your daily activity patterns. Your description should be coherent, utilizing conclusive language to form a well-structured paragraph. Text to Continue: I am $<$INPUT 0$>$ in an urban neighborhood, $<$INPUT 1$>$,…
where $<$INPUT 0$>$ and $<$INPUT 1$>$ are replaced by the candidate personas. $<$INPUT 2$>$ and $<$INPUT 3$>$ will be replaced by the activity habits extracted from historical data. Specifically, we formulate the $<$INPUT 2$>$ in the following format:
$<$INPUT 2$>$ format
During $<$weekend or weekday$>$, you usually travel over $<$distance$>$ kilometers a day, you usually begin your daily trip at $<$time of first daily activity$>$ and end your daily trip at $<$time of last daily activity$>$. you usually visit $<$location of first daily activity$>$ at the beginning of the day and go to $<$location of last daily activity$>$ before returning home.
Pattern evaluation Now you act as a person with the follwing feature: $<$INPUT 0$>$ Here is a plan for you. $<$INPUT 1$>$ On the scale of 1 to 10, where 1 is the least possibility and 10 is the highest possibility, rate how likely you will follow this plan according to your routine pattern and explain the reason. Rating: $<$fill in$>$ Reason:
where $<$INPUT 0$>$ is the candidate pattern. $<$INPUT 1$>$ is the daily activities plan for evaluation.
Evolving-based motivation reterieval Context: Act as a person in an urban neighborhood and describe the motivation for your activities. $<$INPUT 0$>$ In the last $<$INPUT 1$>$ days you have the following activities: $<$INPUT 2$>$ Instructions: Describe in one sentence your future motivation today after these activities. Highlight any personal interests and needs.
where $<$INPUT 0$>$ is replaced by the selected pattern. $<$INPUT 1$>$ is replaced by the number of days from the date to plan activities. $<$INPUT 2$>$ is the historical activities corresponding to the chosen date.
Learning-based motivation reterieval Context: Act as a person in an urban neighborhood. $<$INPUT 0$>$ If you have the following plans: $<$INPUT 1$>$ Instructions: Try to summarize in one sentence what generally motivates you for these plans. Highlight any personal interests and needs.
where $<$INPUT 0$>$ is replaced by the selected pattern. $<$INPUT 1$>$ Is replaced by the retrieved historical activities.
[…]
where $<$INPUT 0$>$ is replaced by the selected pattern. $<$INPUT 1$>$ Is replaced by the retrieved motivation. $<$INPUT 2$>$ Is replaced by the most frequently visited locations.
Pandemic
Now it is the pandemic period. The government has asked residents to postpone travel and events and to telecommute as much as possible.
Appendix B Learning-Based Motivation Retrieval
For the learning-based motivation retrieval, the score approximator is parameterized using a fully connected neural network with the following architecture:
Table 6: Architecture of the score approximator.
Layer Input Size Output Size Notes
Input Layer 3 64 Linear
Activation - - ReLU
Output Layer 64 1 Linear
We include the day of the year for the query date, whether it shares the same weekday as the reference date, and whether both the query and reference dates fall within the same month as input features. Settings for the learning process are as follows: Adam is used as the optimizer, batch size is 64, learning rate is 0.002, and the number of negative samples is 2.